Please answer the questions and provide the graphs.
Antibody & Immunohistochemical Methods In a study of the proliferative activity of breast cancers, Veronese and Gambacorta (A-1) used the Immunohistochemical methods. The investigators obtained tumor tissues from 203 patients with breast ranged in age from 26-82 years. The following table shows the K3-67 values (expressed as percentages) for thes 10 12 10.80 1054 27 30 838 19 30 16.40 440 26 80 16 60 21.42 25 11 1260 1796 4112 2109 11.95 33.30 4.53 1940 15 5.48 23.50 31 4.65 7.84 10.90 9.03 51206.40 13.72 32 90 9.80 243 200 4.099.206.20 1760 1.50.00 | 1000 | 20.00 | 30,00 5.00 15.00 25.00 10.00 12.70 | 3000i 1000 150012000 739 | 400 | 2500 | 2000 | 30.00 21.36 49.85 29.70 1995 5.00 11.36 | 2489 | 2955 | 1000 | 3890 8.12 28.85 1980 499 6.00 9.88 3.14 5.,00 4420 30.00 9 4.33 9.20 487 10.00 29.10 3.00 2.00 8.10 4.84 4.23 10.00 19.83 20.00 4.77 13.11 750 2000 500 455 607 15.00 4000 18.79 13.76 4582 432 569 142 18 57 12.82 4.50 441 1.88 14.24 9.11 5.00 10.00 4.12 414 2.03 2.69 3.69 542 59 10.00 6276.37 13.78 27.559.83 6.55 8.21 342 7.53 3.51 8.58 5.00 2950 9.608.03 1370 | 5.60 | 28.10ェ548-17.00 75 SOURCE SINo M Veronese, Ph.D. Used with permasion (a)Use these data to prepare: A frequency distribution (b) What percentage of the measurements are less than 102 al What percentage of the measurements are between 20 and 49 inclusive? (e) How many of the measurements are greater than 392 What proportion of the subjects have measurements greater than or equal to 20 A relative frequency distributon tve redueney uencybnameone picks a measurement at random from this data set and of the measurements are either less than 10 or greater than 69 What proportion data set and asks you A histogram A frequency polygon guess the value. What would be your answer? Why?
Answer #1

(a)

Class Frequency Relative Frequency Cumulative Frequency Cumulative Relative Frequency
0-9.99 102 0.502 102 0.502
10-19.99 52 0.256 154 0.759
20-29.99 25 0.123 179 0.882
30-39.99 12 0.059 191 0.941
40-49.99 7 0.034 198 0.975
50-59.99 3 0.015 201 0.990
60-69.99 0 0.000 201 0.990
70-79.99 2 0.010 203 1.000
Total 203 1

Histogram :

120 100 80 260 40 20 102 25 12 (0 0 10 20 30 4050 60 70 80

Frequency Polygon :

120 100 80- 2 60 40 20- 0 10 20 30 4050 60 70 80

(b)

Percentage of measurements less than 10 = 102 / 203 = 0.5025 = 50.25%

(c)

Proportion of subjects have measurements greater than equal to 20 = (203 - 154) / 203 = 49 / 203 = 0.2414

(d)

Percentage of measurements between 20 and 49 inclusive = 43 / 203 = 0.2118 = 21.18 %

(e)

Number of measurements greater than 39 = 7 + 3 + 0 + 2 = 12

(f)

Proportion of  measurements either less than 10 or greater than 69 = (102 + 2 ) /203 = 0.5123

(g)

Answer would be 5 because 5 is most frequently occurring observation (Mode)

So, 5 has the highest probability of occurring.